### v2

# EgoOops: A Dataset for Mistake Action Detection from Egocentric Videos referring to Procedural Texts

Abstract
Mistake action detection is crucial for developing intelligent archives that detect workers’ errors and provide feedback. Existing studies have focused on visually apparent mistakes in free-style activities, resulting in video-only approaches to mistake detection. However, in text-following activities, models cannot determine the correctness of some actions without referring to the texts. Additionally, current mistake datasets rarely use procedural texts for video recording except for cooking. To fill these gaps, this paper proposes the EgoOops dataset, where egocentric videos record erroneous activities when following procedural texts across diverse domains. It features three types of annotations: video-text alignment, mistake labels, and descriptions for mistakes. We also propose a mistake detection approach, combining video-text alignment and mistake label classification to leverage the texts. Our experimental results show that incorporating procedural texts is essential for mistake detection. Data is available through https://y-haneji.github.io/EgoOops-project-page/.
Fig. 1: An example task (cardboard) from our EgoOops dataset. EgoOops includes 50 egocentric videos across five procedural domains and corresponding procedural texts. It contains three types of annotations: video-text alignment, mistake labels, and descriptions explaining the errors in each segment.
Index Terms—  Mistake detection, Procedural activity, Video dataset, Egocentric video
1 Introduction
Table 1: Comparison of mistake datasets. Range: mistake labels to a video or each step segment with start and end times; Cat.: finer-grained categorization than correct , mistake , and correction ; Desc.: descriptions explaining why each segment is incorrect; OM, EM: order mistakes, execution mistakes (see Section 3.1); Proc.: workers follow step-by-step procedural texts. Domain specific (e.g., extra screws). Orders unseen in the training set, not always faulty.
Dataset Mistake annotations Domain Proc. Ego #videos Duration (hour)
Range Cat. Desc. OM EM
Assembly101 [1] Segment $\times$ $\times$ $\checkmark$ $\times$ Assembly $\times$ $\checkmark$ 1,425 167
ATA [2] Video $\checkmark^{\ast}$$\times$ $\checkmark^{\dagger}$checkmark Assembly $\times$ $\times$ 1,152 24.8
HoloAssist [3] Segment $\times$ $\checkmark$ $\times$ $\checkmark$ Assembly $\times$ $\checkmark$ 2,221 166
IndustReal [4] Segment $\times$ $\checkmark$ $\checkmark$ $\checkmark$ Assembly $\times$ $\checkmark$ 84 5.8
EgoPER [5] Segment $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Cooking $\checkmark$ $\checkmark$ 386 28
CaptainCook4D [6] Segment $\checkmark^{\ast}$checkmark $\checkmark$ $\checkmark$ Cooking $\checkmark$ $\checkmark$ 384 94.5
EgoOops (Ours) Segment $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Diverse $\checkmark$ $\checkmark$ 50 6.8
Procedural activities are common in daily life and expert fields, such as assembly, experimentation, and cooking. People often carry them out by following procedural texts in the real world. During this process, mistakes negatively impact quality, speed, cost, and safety. Common mistakes include skipped necessary steps or wrong execution ways, which can sometimes result in life-or-death situations. One promising solution to this problem is to develop an intelligent video archive that records workers’ activities, detects their mistakes, and shows them the mistake clips to prevent recurrences.
The intelligent video archives are developed using video datasets recording workers’ steps in detail. Many egocentric video datasets [7, 8, 9, 10, 1, 11, 3, 6, 4, 5] have been proposed by equipping a worker with an egocentric camera to capture activity details. Previously, most datasets were interested in correct execution of activities [7, 8, 10, 9, 11]. Some recent studies have also included and annotated mistake actions in their assembly [1, 3, 2, 4] and cooking [6, 5] video datasets. Using these mistake video datasets, researchers have proposed mistake detection approaches [1, 2, 6, 12, 13, 5].
However, these studies have the following three limitations (see Table 1). L1: video-focused approaches.Existing studies [1, 2, 6, 12, 13, 5] have focused on video-only approaches to mistake detection and not utilized procedural texts. Their approaches are suitable for visually apparent mistakes like incorrectly attached parts and falling plates. Nevertheless, in text-following activities, some mistakes are deviations from procedural texts, thus not obvious only from visual cues (e.g., in Fig. 1, the use of tape is a mistake because the text designates glue). Therefore, besides videos, texts are essential for models to detect mistakes accurately. L2: specific domains.Procedural texts are included in only a few mistake datasets recording cooking [6, 5]. Since many real-world activities follow procedural texts, it is essential to collect data from more domains. L3: rough mistake labels.Most datasets define coarse-grained (correct /mistake /correction[1, 3, 4] or domain-specific (e.g., extra screws[2, 6] categories. General and fine-grained categorization enables analysis of mistake patterns across diverse domains (e.g., a commonly frequent category of mistakes).
To address these issues, we propose a novel dataset called EgoOops (see Fig. 1), where egocentric videos record erroneous activities when following procedural texts (L1) across diverse domains (L2). Given the collected videos and texts, we perform the following three steps for annotations. First, we align steps in the procedural text with video segments (i.e., start and end timestamps). Second, if a segment contains a mistake action, we categorize it into six mistake classes (L3). Finally, the segments assigned mistake labels are provided with descriptions of why the actions are considered mistakes. As for the size and tasks, EgoOops contains 50 egocentric videos totaling 6.8 hours across five tasks of new domains: electrical circuits, color mixture experiments, ionic reaction experiments, toy building blocks, and cardboard crafts.
We also propose an approach to the problem of mistake action detection, especially focusing on the utilization of procedural texts (L1). To leverage the texts, our approach combines video-text alignment and mistake label classification; the former localizes the start and end times of each procedural step, and the latter assigns the step segments one label of mistake classes. In experiments using EgoOops, our multi-modal approach outperforms a video-only baseline, and the ablation of textual inputs decreases our performance. These results demonstrate that incorporating procedural texts is essential for mistake detection.
2 Related work
Mistakes are crucial issues in procedural activities, leading to recent proposals of various datasets (see Table 1). These are categorized into free-style and text-following settings.
In free-style activity datasets [1, 2, 3, 4], workers aim to complete goals not relying on procedural texts. For example, Assembly101 [1] records toy assembly and annotates segment-level labels of correct , mistake , and correction . Building on such datasets, previous studies have proposed various approaches to find mistakes in videos [1, 2, 3, 12, 13]. These approaches rely only on videos because mistakes in free-style activities are mainly visually apparent such as incorrectly attached parts and dropped screws.
Another line of studies records workers following procedural texts in their datasets [6, 5]. EgoPER [5] annotates recipe-execution videos with order and execution mistakes, and CaptainCook4D [6] proposes categorization specific to cooking (e.g., temperature error ). One problem is that they still do not utilize procedural texts to find mistakes in videos. In text-following activities, some mistakes are deviations from procedural texts, thus the texts are essential for models. In addition, to the best of our knowledge, no datasets besides cooking involve procedural texts, whereas many real-world activities follow them. Our EgoOops dataset annotates video-text alignment and mistake labels across diverse domains, promoting the utilization of texts in mistake detection.
3 EgoOops dataset
Our dataset consists of procedural texts and egocentric videos containing mistake actions. This section describes terms, text preparation, recording process, annotations, and statistics.
3.1 Mistake definition
We define mistakes as deviations from a procedural text. This definition leads to two mistake types: order mistakes and execution mistakes . Order mistakes occur when the steps executed by a worker are different from the ones listed in the procedural text. These include skipping necessary steps (missing ), swapping the step order (out-of-order ), inserting extra steps (undefined ), and pausing a step and resuming after others (split ). Execution mistakes occur when a worker fails to follow the instructions while carrying out a step. We categorize them into the following six classes: 1. working with wrong objects (object ), 2. grasping wrong objects and releasing them without using (mispick ), 3. correction of mistake actions (correction ), 4. unintended actions (accident ), 5. performing in the wrong way (way ), 6. others (others ).
3.2 Task selection and text preparation
To record procedural activities in diverse domains with various mistake actions, we selected the following five tasks: electrical circuits (EC ), color mixture experiments (CM ), ionic reaction experiments (IR ), toy building blocks (BB ), and cardboard crafts (CB ). Their domains are diverse (i.e., electronics, chemistry, assembly, and crafts), and various mistakes can happen (e.g., using wrong chemicals, accidentally cutting off cardboard).
One of the authors wrote a procedural text for each task before recording. They referred to the web for color mixture and cardboard crafts, and texts for ionic reactions and circuits are borrowed from out-of-the-box kits. For building blocks, they wrote it from scratch because we found no references. Note that all texts were originally written in Japanese, then we manually translated them to prepare an English version.
3.3 Video recording
Four Japanese graduate students (4 males) performed the tasks following the texts while equipped with a head-mounted camera. A participant worked on every task two or four times, totaling 10 recordings for each task. To record various mistake actions, they intentionally performed mistake actions that we assumed in advance (e.g., skipping several steps). Out of ten times to record videos, they contained the intentional mistakes five times and tried to follow procedures correctly for the other five times. Note that accidental mistakes due to carelessness also happened even in the latter.
3.4 Annotations
EgoOops contains annotations of video-text alignment, mistake labels, and their descriptions. One person annotates the whole dataset using a web annotation tool described in the supplementary materials.
We first annotate video-text alignment by extracting start and end timestamps (i.e., segments) and mapping them to the corresponding steps in a procedural text. Note that we also annotate steps not written in the procedural text as undefined (e.g., grasping the wrong object). The video-text alignment itself reveals order mistakes. For example, if step 2 occurs before step 1, these steps are considered swapped. Thus, we do not explicitly annotate order mistakes.
If a segment indicates an execution mistake, we assign one label of the six execution mistake classes listed in Section 3.1. This categorization can be used to analyze mistake patterns across diverse tasks (e.g., a commonly frequent category of mistakes). In addition, we also provide the mistake segments with descriptions of why the performed actions are considered mistakes. They can be used to analyze model performance in detail or interact with workers through smart assistants.
We ensure the quality of the original annotations by testing for agreement with another annotator. For video-text alignment, the annotator newly extracts segments and maps them to the corresponding step labels, resulting in the temporal Intersection over the Union (tIoU) to the original at 88.8. In addition, they watch the original segments and give new mistake labels and descriptions. The labels and descriptions are compared using Cohen’s kappa [14] and BERTScore [15], achieving 86.8 and 96.3, respectively. These high scores confirm the quality of the original annotations.
3.5 Statistics
In this section, we first report video- and text-side statistics on EgoOops, then discuss mistake label statistics. The statistics of mistake descriptions are written in the supplementary materials.
Table 2 shows different trends between the tasks in terms of video duration, segment duration, and the number of segments. For video duration, the longest is the cardboard task at 26.1 minutes, while the shortest is the building block task at 1.9 minutes. As for the texts, Table 2 compares the number of steps in a procedural text and the number of words per step. The task with the most steps is the cardboard task, while the task with the longest instructions is the building block task. These statistics indicate that EgoOops covers a variety of procedural tasks ranging from short to long.
Table 2: Statistics of recorded videos and procedural texts.
Task Videos Segments Texts
#vid. Avg. (min) #seg. #seg. / #vid. Avg. (sec) #steps per text #words / #steps
EC 10 3.2 98 9.8 15.4 8 7.6
CM 10 4.4 91 9.1 25.8 8 17.0
IR 10 5.4 95 9.5 29.7 9 12.7
BB 10 1.9 87 8.7 9.0 7 18.6
CB 10 26.1 167 16.7 86.7 14 9.6
All 50 8.2 538 10.8 40.7 9.2 13.5
Table 3 shows comparisons of execution and order mistakes between tasks. Of all the mistakes, 41% ($=95/233$) are execution mistakes, and the others are order mistakes. Among the order mistakes, splitting steps is the most frequent. Table 4 shows the counts of the labels for execution mistakes. The two most frequent labels are grasping the wrong objects and releasing them without use (label 2) and working in the wrong way or moving (label 5). In addition, we also find unique mistake patterns of each task (e.g., unintended actions (label 4) frequently happen in ionic reaction experiments but rarely in the other tasks).
Table 3: Comparisons of execution and order mistakes.
Task Execution mistake Order mistake
Missing Out-of-order Split Undefined
EC 22 2 10 12 6
CM 22 1 2 3 9
IR 19 2 4 1 6
BB 19 0 3 6 10
CB 13 7 21 29 4
Total 95 12 40 51 35
Table 4: Statistics of mistake labels.
Task Mistake label
1. object 2. mispick 3. correction 4. accident 5. way 6. others
EC 9 5 1 2 3 2
CM 4 8 0 2 5 3
IR 0 3 1 5 6 4
BB 2 5 5 1 5 1
CB 5 3 0 1 2 2
Total 20 24 7 11 21 12
4 Application: mistake action detection
We propose a text-oriented approach to mistake action detection. This section formalizes the problem, explains our approach, and reports experimental results on EgoOops.
4.1 Problem formalization
Mistake action detection consists of two problems: video-text alignment and mistake label classification. It finally localizes temporal segments and classes of mistakes in a video.
Given an untrimmed video $\mathbf{V}=(\mathbf{f}_{1},\dots,\mathbf{f}_{L})$ and the corresponding procedural text $\mathbf{T}=(\mathbf{t}_{1},\dots,\mathbf{t}_{K})$, our first problem is video-text alignment. Here, $\mathbf{V}$ consists of $L$ frames, and $\mathbf{T}$ includes $K$ steps of instructions. For $k$-th step $\mathbf{t}_{k}$, our goal is to localize the start and end frame numbers $(s_{k},e_{k})$.
The alignment outputs are passed to the next problem of mistake label classification, which predicts a mistake label for each step segment. The $k$-th step segment is represented by extracted video frames based on the start and end frame numbers as $\mathbf{V}^{\prime}_{k}=(\mathbf{f}_{s_{k}},\dots,\mathbf{f}_{s_{e}})$. Our objective is to assign the segment $\mathbf{V}^{\prime}_{k}$ one label of mistake classes. In our experiments, we group the mistakes except for 3. correction (see Section 3.1) into the common mistake class and solve the classification of three classes: correct , mistake , and correction .  [^1] Our preliminary experiments showed that distinguishing the seven classes (six classes in Section 3.1 plus the correct class) is difficult currently.
4.2 Models
Video-text alignment. We adopt StepFormer [16] to tackle video-text alignment. Stepformer was originally proposed to learn video-text alignment from untrimmed videos with narrations in a self-supervised fashion [16]. We train it to fit EgoOops better in the following two steps. First, we apply its self-supervised learning to pre-train it on Ego4D [10] using EgoVLPv2 [17] features. Second, we add a fully-supervised loss to learn from video-text alignment annotations and fine-tune the pre-trained model on EgoOops. Specifically, we improve the similarity between the remaining slots [16] and annotated video segments, and vice versa, using InfoNCE contrastive loss [18]. We call the resulting model StepFormer++ to distinguish it from the original one.
Mistake label classification. We train a multi-modal classifier for mistake label classification. First, the trimmed video clip and step instruction are separately converted into features using EgoVLPv2 [17]. Then, the video features are averaged along the temporal dimension and concatenated to the text features. Finally, the concatenated vectors are forwarded to a two-layer perceptron to predict the logits of the mistake classes. Regarding model training, we compute the class-balanced focal loss [19] because of the relatively low frequency of the mistake and correction classes. We feed the ground-truth step segments to stabilize training whereas the predicted ones are used for testing [20, 21, 22].
4.3 Experiments on mistake action detection
Baseline. We compare our approach with a recent temporal action localization (TAL) model ActionFormer [23] because of TAL’s similarity to our problem settings. [^2] Existing mistake action detection approaches do not fit our settings. For example, the closest one is EgoPED as it predicts both temporal segments and mistake labels [5]. However, its anomaly detection approach predicts binary labels of correct or mistake , thus it cannot solve our three class settings. In our experiments, we train ActionFormer to predict mistake labels for the detected segments, instead of action labels as in TAL.
Evaluation metrics. We follow TAL [23] and report mean average precision (mAP) at tIoU thresholds of 0.1, 0.2, and 0.3. It computes the mean of average precision across only mistake and correction classes because we focus on mistake detection performance. Since our problem formulation involves video-text alignment, our metric requires correctly predicting both step and mistake labels. Note that ActionFormer processes only videos and does not conduct video-text alignment [23]; yet we assign step labels to the segments sequentially from the video’s start to end for fair comparison.
Results. Our proposed method achieves an average mAP of 2.5, surpassing ActionFormer’s 0.7 (see Table 5). In contrast, our score is still far from the oracle’s 34.7, which uses ground-truth step segments and only performs mistake label classification. This suggests that accurate video-text alignment largely improves mistake action detection. In addition, we explore the success and failure examples visually (see Fig. 2). In the shown success case, the correct alignment leads to the finding of a mistake; in the shown failure case, the localized step is incorrect, hurting the performance of mistake classification. Moreover, we ablate the text inputs into our classifier and allow it to use only videos, then our average mAP decreases to 0.3 (see Table 6). This demonstrates the necessity of text information for accurate mistake classification.
Table 5: Results of mistake action detection. Oracle denotes upper-bound performance using ground-truth step segments.
Methods mAP@tIoU
0.1 0.2 0.3 Avg.
ActionFormer [23] 1.8 0.1 0.1 0.7
StepFormer++ w/ MLP (ours) 2.6 2.5 2.5 2.5
GT steps (oracle) 34.7
Fig. 2: Success (left) and failure (right) cases of mistake detection. Text boxes at the bottom show the predicted steps.
Table 6: Ablation study of textual inputs to the mistake label classifier of our approach.
Inputs to classifier Avg. mAP Avg. mAP (oracle)
Video Text
$\checkmark$ 0.3 4.7
$\checkmark$ $\checkmark$ 2.5 34.7
4.4 Experiments on video-text alignment
We compare our StepFormer++ with two versions of the original StepFormer [16] by evaluating video-text alignment performance. One version is trained on Ego4D and evaluated in a zero-shot manner (ZS), and the other is further fine-tuned on EgoOops in a self-supervised manner (SS). In Table 7, we report frame-wise F1, Precision, Recall, and Mean over Frames (MoF) [16]. StepFormer++ achieves an F1 score of 28.1 outperforming both ZS and SS of 24.1 and 26.1, respectively. The results confirm that our fully supervised loss trains StepFormer well if alignment annotations are available.
Table 7: Results of video-text alignment. ZS: zero-shot, SS: self-supervised, FS: fully supervised.
Methods F1 Prec. Rec. MoF
StepFormer (ZS) [16] 24.1 24.6 23.7 24.9
StepFormer (SS) [16] 26.1 26.6 25.7 27.0
StepFormer++ (ours, FS) 28.1 28.4 27.9 28.1
4.5 Experiments on mistake label classification
Classifiers. EgoOops dataset contains tasks of new, various domains (see Section 3.2) other than existing assembly [1, 2, 3, 4] and cooking [5, 6] datasets. To confirm the novelty of our dataset, we test classification models trained on such other datasets. Specifically, we compare our classifier trained on EgoOops with ones proposed by and trained on Assembly101 [1] and CaptainCook4D [6].
MLLMs. Finding mistake actions is a visual reasoning task; models understand the video and procedural text and decide whether a worker follows the step instructions correctly. Multi-modal large language models (MLLMs) perform well in visual reasoning benchmarks [24, 25], thus we instruct them to solve mistake label classification given a trimmed video clip, the task’s procedure, and the performed step. Our experiments adopt leading open-source models: InternVL2.5-8B [24] and Qwen2-VL-7B-Instruct [25].
Results of classifiers. Table 8 suggests that our new domains are challenging for the existing models; our classifier achieves an F1 score of 41.3 for the mistake class, but Assembly101 and CaptainCook4D record 14.8 and 27.1, respectively.
Results of MLLMs. MLLMs find mistake samples accurately but struggle with correct samples (see Table 8). This suggests their limited capability for reasoning.
Table 8: Results of mistake action classification. ZS: zero-shot, FS: fully-supervised. Note that [6] addresses binary classification of correct or mistake , thus the scores for correction are not available.
Methods Mistake Correction
Prec. Rec. F1 Prec. Rec. F1
Classifier Uniform sampling 16.4 33.3 21.9 1.3 33.3 2.5
Assembly101 [1] (ZS) 14.8 14.8 14.8 2.4 42.9 4.5
CaptainCook4D [6] (ZS) 16.5 76.1 27.1 N/A
MLP (ours, FS) 35.0 48.9 40.8 57.1 57.1 57.1
MLLM InternVL2.5-8B [24] 47.6 34.1 39.7 0.0 0.0 0.0
Qwen2-VL-7B-Instruct [25] 75.0 30.7 43.5 2.6 14.3 4.4
5 Conclusion
This paper proposed the EgoOops dataset, which consists of egocentric videos, procedural texts, and three types of annotations: video-text alignment, mistake labels, and descriptions. We also proposed a text-oriented approach to the problem of mistake action detection. Our experiments on EgoOops revealed that text information is essential for mistake detection.